OpenStudy (anonymous):

Write an equation of each line in standard form with integer coefficients. 13. y=3/2x-1/2 14. y=3/2x-1/4 15. y = 4.2x + 1.8 16.

OpenStudy (***[isuru]***):

hi, first one... y = (3/2)x - 1/2 multiply the whole by 2 2y = 3x -1 second one... y = (3/2)x -1/4 multiply the whole by 4 4y = 6x - 1 third one.... y = 4.2x + 1.8 multiply the whole by 10 10y = 42x +18 hope this will help ya!!!!

OpenStudy (anonymous):

is the last one 42x+18 or 4.2x+1.8